The fishing pole shown makes an angle of 20.0° with the horizontal. What is the torque exerted by the fish about an axis perpendicular to the page and passing through the angler’s hand if the fish pulls on the fishing line with a force F = 100 N at an angle 37.0° below the horizontal? The force is applied at a point 2.00 m from the angler’s hands. 


1
Expert's answer
2021-04-19T17:11:34-0400


Torque is the twisting force acting on a body. The torque acting on a body is given as:

τ = r Fsinθ

r = radius

F = acting force

The torque acting perpendicular to the paper is represented below:



The total angle at which the force acting is given as:

θ = 20º+37º = 57º

r = 2 m

F = 100 N

"τ = 2 \times 100 \times sin57º \\

= 200 \times 0.838 \\

= 167.7 \;N \times m"

Answer: 167.7 Nm